OpenStudy (anonymous):

How many ohms of resistance are in a 120-volt hair dryer that draws 7.6 amps of current?

OpenStudy (shane_b):

Use Ohm's law:\[V=IR\]You know V and I...just solve for R.

OpenStudy (anonymous):

R=V/I so, you have R= 120/7.6 R= 15.79 ohms

OpenStudy (anonymous):

You have to write the equation in function of Resistance.

OpenStudy (anonymous):

just use ohm's law: r=v/i

OpenStudy (shane_b):

I would hope that after 8 months the original poster was way past this question :)

OpenStudy (anonymous):

thank you !